cos(x+pi/2) = -sin(x)
Answer by MathLover1(20855)


cos(x+pi/2)=-sin (x)

Solution:
Transform the left side of the identity until is is the same as the right side.
cos(x + pi/2) = -sin (x)
cos(pi/2 - (-x)) = -sin (x)

Apply the co-function property:
 [cos(pi/2 + x) = sin (x)]
sin (-x) = -sin (x)

Apply the even/odd formula:
[sin (-x) = - sin (x)]
-sin (x) = - sin (x)